Circular No. 6871

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998an IN UGC 3683 Ron Arbour, South Wonston, Hants, England, reports his discovery of

an apparent supernova (mag ∼ 15.3) on an unfiltered CCD image obtained with a 0.30-m f/4.5 reflector on Apr. 6.871 UT. The new star is located at α = 7^{h}08^{m}17^{s}.04, δ = +46^{o}06′55′′.3 (equinox 2000.0), which is ∼ 32′′ east and 3′′ south of the center of UGC 3683. T. Boles, Wellingborough, confirmed the star on two exposures taken at about the same time, and Arbour obtained a total of twelve images on the discovery night and another image on the following night. The object was not visible during Arbour’s previous observation of the same galaxy on Mar. 30, nor does it appear on the Digitized Sky Survey.

CI CAMELOPARDALIS R. I. Hynes, P. Roche, and C. A. Haswell, University of Sussex; J.

Telting, Isaac Newton Group; and M. Lehnert and Y. Simis, University of Leiden, report: “UBVRI photometry of CI Cam was obtained with the Jacobus Kapteyn Telescope at La Palma on Apr. 3.87 UT. We find U = 10.3, B = 10.7, V = 9.7, R = 8.6, I = 7.9, with errors of ∼ 0.1 mag. The source had therefore faded by 0.5 mag in B and V since the observation 18 hr earlier of Garcia et al. (IAUC 6865). Our colors are very similar to those of Bergner et al. (1995, A.Ap. Suppl. 112, 221), indicating that the outburst optical spectral energy distribution is the same as that in quiescence, but brighter. Assuming visual extinction of 1.5 mag (Chkhikvadze 1970, Astrofizika 6, 65), the intrisic flux distribution is approximately a power-law, with a power-law index of −1.2.”

IRAS 18325–5926 A. C. Fabian, J. C. Lee, and K. Iwasawa, Institute of Astronomy, Cam-

bridge; K. Jahoda, Goddard Space Flight Center; W. N. Brandt, Pennsylvania State University; and C. S. Reynolds, JILA, University of Colorado, report: “The 11-hr periodicity found in the power spectrum of the RXTE x-ray lightcurve of the active galaxy IRAS 18325–5926 (IAUC 6835) is an artifact produced by the standard background-modeling software. There are therefore no indications of an imminent blackhole merger. The evidence for a 16-hr periodicity of the source in Mar. 1997 (Iwasawa et al. 1998, MNRAS 295, L20) is unaffected, since it was taken with the imaging x-ray satellite ASCA.”

1998 April 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green